OpenStudy (anonymous):

(5/3) + (3/x+3) =5/3

OpenStudy (anonymous):

\[5/3 +3/x+3=5/3\]

OpenStudy (anonymous):

I am just not sure what to do to find x.

OpenStudy (anonymous):

\[\frac{ 5 }{ 3} + \frac{ 3 }{ x+3} = 5/3\] \[=> 5x + 15 + 9 /3x +9 = 5/3\] =>\[5x + 24 /3x + 9 = 5/3\] \[5x + 24 = 14x + 45 /3\] \[=>15x + 72 = 14x + 45\] \[=>x = -72 + 45 = -27\]

OpenStudy (anonymous):

there's a typo error there....

OpenStudy (anonymous):

my answer may be wrong

OpenStudy (anonymous):

it's okay. It helped and I can probably do it now.

OpenStudy (anonymous):

I looked in the answer book and it says 1/(1-3)=0????

OpenStudy (anonymous):

I just don't understand how they got that.

OpenStudy (anonymous):

x=0?

OpenStudy (anonymous):

I don't know. I am really confused by that.